# Linear Complexity Gibbs Sampling for Generalized Labeled Multi-Bernoulli Filtering

### v1

@@ IV-B Tracking Accuracy
The resultant OSPA and OSPA${}^{(2)}$ for varying numbers of trajectories and iterations are shown in Figs. 5(a) and (b). For a relatively small number of objects, the average filtering and tracking errors for all GS variants are virtually identical. As the number of objects is gradually increased, as expected, the corresponding errors for $SGS^{+}$ (the most expensive variant) exhibit the smallest increase of all variants. Conversely but also as expected, the errors for $RGS^{+}$ show the largest increase, due to slow mixing and insufficient diversity. The errors for $TGS^{+}$ are significantly better than those for $RGS^{+}$, even though both have the same linear complexity, due to the combination of tempering with smart coordinate selection. The errors for $\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ are higher than that of $TGS^{+}$ but are generally lower than for $RGS^{+}$. The plots for varying numbers of iterations similarly indicate that $SGS^{+}$ generally has the lowest error, while $RGS^{+}$ has the highest, and $TGS^{+}$/$\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ are comparatively robust. The effect of varying the detection and clutter rate is shown in Figs. 5(c) and (d). Though $SGS^{+}$ generally outperforms $TGS^{+}$/$\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$, the gap appears to narrow as the SNR increases. Again, $TGS^{+}$ outperforms $\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$which are better than $RGS^{+}$ in all cases.
Overall, the results suggest that $TGS^{+}$/$\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ generally outperform $RGS^{+}$ on tracking accuracy, even though they have the same order complexity. In addition, there is small disparity between $\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$, suggesting that scan order can influence results. Unsurprisingly, it is observed that $SGS^{+}$ generally outperforms $TGS^{+}$, due to a significant disparity in order complexity and hence running times. Consequently, the proposed linear complexity $TGS^{+}$ based solutions have the potential to offer a robust trade-off between tracking performance and computational load.
Figure 6: Experiments for sample diversity.
@@ IV-C Sample Diversity
The trend in Fig. 6(a) confirms that the number of unique samples follows the increase with the number of trajectories, which is necessary to capture the corresponding increase in the diversity of the components in the filtering density. In the case of the lowest number of objects, it can be seen from in Fig. 5(a) that even though all the GS variants have the same tracking accuracy, they generally produce different numbers of unique solutions. This is due to the fact that a small number of samples is still sufficient to capture the GLMB filtering density. As the number of objects increases, it can be seen that $SGS^{+}$ produces over 2 times more unique samples than $TGS^{+}$, but incurs a significant increase in time complexity, which is expected due to the exhaustive traversal of all coordinates in $SGS^{+}$. In addition, $TGS^{+}$ produces up to 2 times more unique samples than $RGS^{+}$, even though both have the same complexity, due to the smart (single) coordinate selection in $TGS^{+}$. The number of unique samples from $\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ is also higher than that from $RGS^{+}$ but lower than that from $TGS^{+}$. The relative trends in the number of unique samples are generally consistent with the relative trends in the tracking error in Fig. 5(a).
@@
Examination of Fig. 6(b) confirms that the number of iterations is directly influences to the number of unique samples. It can also be seen that $TGS^{+}$ produces up to 2 times the number of unique samples compared to $RGS^{+}$ with the same order complexity. While $\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ produce a slightly lower number of unique samples than $TGS^{+}$, there is also a corresponding reduction in the measured run times. As expected, $SGS^{+}$ outperforms the other linear complexity variants, but at a significant increase in computational complexity. Again, it is observed in general that the relative increase in the number of unique samples is consistent with the increasing running times in Fig. 3(b) and with the decreasing OSPA and OSPA${}^{(2)}$ values in Fig. 5(b). Inspection of the curves for increasing detection rates in Fig. 6(c) shows a decreasing trend due to lower number of components required to accurately represent the GLMB filtering density in higher SNR. Inspection of the curves for increasing clutter rates in Fig. 6(d) shows the converse trend.
@@
It can be seen that $TGS^{+}$/$\overrightarrow{DGS}$${}^{+}$/$\overleftarrow{DGS}$${}^{+}$ generally outperform $RGS^{+}$ on increased sample diversity, and that the former have the same (or better) order complexity than the latter. Compared to $RGS^{+}$, the experimental results in general suggest that the use of tempering and/or state informed coordinate selection can significantly improve slow mixing and sample diversity, but nonetheless retain the benefit of linear complexity.
@@ V Conclusion
A linear complexity Gibbs Sampling (GS) framework for GLMB filtering density computation has been developed. Specifically, we use the recently developed tempered GS approach to generate significant component of the GLMB filtering density based on measurements received. Our so-called $TGS^{+}$ framework tailors tempered GS to exploit the structure of the problem for scalable GLMB filtering, and offers trade-offs between tracking accuracy, computational efficiency, and memory loads. Further, this framework enables the SGS algorithm in [30] to be implemented with one order of magnitude reduction in complexity, although this is still higher than linear. Comprehensive numerical experiments compare the performance of the linear complexity $TGS^{+}$, and the two special cases of deterministic and completely random-scan GS, namely $DGS^{+}$ and $RGS^{+}$. Our results indicate that, of the linear complexity solutions, $TGS^{+}$ provides the best tracking accuracy. $DGS^{+}$ is simple to implement and slightly faster, with a very small degradation in tracking performance. While $RGS^{+}$ is the cheapest, its tracking performance degrades significantly. Optimizing $TGS^{+}$ could yield better balance between computational load and tracking accuracy, and is a prospective venue for investigation. Due to the effectiveness of the proposed approach, extension to multi-dimensional assignment problem could alleviate the computational bottlenecks in multi-scan and/or multi-sensor problems.

### v2

@@ IV-B Tracking Accuracy
Figure 6: Experimental results for sample diversity.
The resultant OSPA and OSPA${}^{(2)}$ for varying numbers of trajectories and iterations are shown in Figs. 5(a) and (b). For a relatively small number of objects, the average filtering and tracking errors for all GS variants are virtually identical. As the number of objects is gradually increased, as expected, the corresponding errors for $SGS^{+}$ (the most expensive variant) exhibit the smallest increase of all variants. Conversely but also as expected, the errors for $RGS^{+}$ show the largest increase, due to slow mixing and insufficient diversity. The errors for $TGS^{+}$ are significantly better than those for $RGS^{+}$, even though both have the same linear complexity, due to the combination of tempering with smart coordinate selection. The errors for $\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ are higher than that of $TGS^{+}$ but are generally lower than for $RGS^{+}$. The plots for varying numbers of iterations similarly indicate that $SGS^{+}$ generally has the lowest error, while $RGS^{+}$ has the highest, and $TGS^{+}/\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ are comparatively robust. The effect of varying the detection and clutter rate is shown in Figs. 5(c) and (d). Though $SGS^{+}$ generally outperforms $TGS^{+}/\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$, the gap appears to narrow as the SNR increases. Again, $TGS^{+}$ outperforms $\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ which are better than $RGS^{+}$ in all cases.
Overall, the results suggest that $TGS^{+}/\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ generally outperform $RGS^{+}$ on tracking accuracy, even though they have similar complexities. In addition, there is small disparity between $\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$, suggesting that scan order can influence results. Unsurprisingly, it is observed that $SGS^{+}$ generally outperforms $TGS^{+}$, due to a significant disparity in complexity and hence running times. Consequently, the proposed linear complexity $TGS^{+}$ based solutions have the potential to offer a robust trade-off between tracking performance and computational load.
@@ IV-C Sample Diversity
The trend in Fig. 6(a) confirms that the number of unique samples follows the increase with the number of trajectories, which is necessary to capture the corresponding increase in the diversity of the components in the filtering density. In the case of the lowest number of objects, it can be seen in Fig. 5(a) that even though all the GS variants have the same tracking accuracy, they generally produce different numbers of unique solutions. This is due to the fact that a small number of samples is still sufficient to capture the GLMB filtering density. As the number of objects increases, it can be seen that $SGS^{+}$ produces more than twice the number of unique samples of $TGS^{+}$, but incurs a significant increase in time complexity, which is expected due to the exhaustive traversal of all coordinates in $SGS^{+}$. In addition, the informed (single) coordinate selection in $TGS^{+}$ produces more unique samples than $RGS^{+}$ even though both have the same complexity. Note also from Fig. 6(a) that the number of unique samples from $\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ is also higher than that from $RGS^{+}$ but lower than that from $TGS^{+}$. The relative trends in the number of unique samples are generally consistent with those of the tracking error shown in Fig. 5(a).
@@
Examination of Fig. 6(b) confirms that the number of iterations directly influences to the number of unique samples. It can also be seen that $TGS^{+}$ produces up to twice the number of unique samples as $RGS^{+}$. While $\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ produce slightly fewer unique samples than $TGS^{+}$, there is also a corresponding reduction in the observed run times. As expected, $SGS^{+}$ outperforms the other linear complexity variants, but at a significant increase in computational complexity. Again, it is observed in general that the relative increase in the number of unique samples is consistent with the increasing running times in Fig. 4(b) and with the decreasing OSPA and OSPA${}^{(2)}$ values in Fig. 5(b). Inspection of the curves for increasing detection rates in Fig. 6(c) shows a decreasing trend due to lower number of components required to accurately represent the GLMB filtering density in higher SNR. Inspection of the curves for increasing clutter rates in Fig. 6(d) shows the converse trend.
@@
It can be seen that $TGS^{+}/\overrightarrow{DGS}$/${}^{+}/\overleftarrow{DGS}$${}^{+}$ generally outperform $RGS^{+}$ on increased sample diversity, and that the former has the same (or smaller) complexity than the latter. Compared to $RGS^{+}$, the experimental results in general suggest that the use of tempering and/or state informed coordinate selection can significantly improve slow mixing and sample diversity, but nonetheless retain the benefit of linear complexity. Furthermore, the improvement in tracking accuracy of each method levels off once enough significant samples have accumulated.
Figure 7: Experimental results for mixture rate and tempering level.
IV-D Mixture Rate and Tempering Level
Fig. 7 shows the experimental results for $TGS^{+}$, in 2D histograms, with varying mixture and tempering parameters, $\alpha$ and $\beta$. Recall from (21) that higher values $\alpha$ or $\beta$ weaken the effect of tempering. Note from the averaged wall-clock times over all scenarios shown in Fig. 7(a) that weak tempering or mixing slightly reduces computation time. More unique solutions require slightly longer run times due to the removal of duplicated samples. Tracking accuracy (in OSPA${}^{(2)}$) generally improves with stronger tempering or mixing (low values of $\alpha$ or $\beta$) as shown in Fig. 7(b). However, tracking performance may degrade with excessive tempering that over flattens the significant modes of the original distribution. From the average number of unique samples shown in Fig. 7(c), observe that stronger tempering or mixing can increase sample diversity, although excessive tempering can reduce sample diversity for a low memory budget implementation since the samples are no longer a good representation of the stationary distributions. It is, however, difficult to determine the optimal choices of $\alpha$ and $\beta$ because they are heavily scenario-dependent.
@@ V Conclusion
A linear complexity Gibbs Sampling (GS) framework for GLMB filtering density computation has been developed. Specifically, we use the recently developed tempered GS approach to generate significant component of the GLMB filtering density based on measurements received. Our so-called $TGS^{+}$ framework tailors tempered GS to exploit the structure of the problem for scalable GLMB filtering, and offers trade-offs between tracking accuracy, computational efficiency, and memory load. Further, this framework enables the SGS algorithm in [30] to be implemented with one order of magnitude reduction in complexity, although this is still higher than linear. Comprehensive numerical experiments compare the performance of the linear complexity $TGS^{+}$, and the two special cases of deterministic and completely random-scan GS, namely $DGS^{+}$ and $RGS^{+}$. Our results indicate that, of the linear complexity solutions, $TGS^{+}$ provides the best tracking accuracy. $DGS^{+}$ is simple to implement and slightly faster, with a very small degradation in tracking performance. While $RGS^{+}$ is the cheapest, its tracking performance degrades significantly. Optimizing $TGS^{+}$ could yield better balance between computational load and tracking accuracy, and is a prospective venue for investigation.
Due to the effectiveness of the proposed approach, extension to the multi-dimensional assignment problem [37] could alleviate the computational bottlenecks in multi-scan and/or multi-sensor truncation. The recent multi-sensor multi-scan GLMB smoother proposed in [12] extends the SGS solution to the multi-dimensional assignment problem. While this is the first solution to address multi-dimension assignment problems of such large scale, we envisage that its time complexity can be drastically reduced using our proposed TGS approach.
VI Appendix
VI-A Proof of Proposition 2
Proof.
  1. (i)
    For $j<1$, it follows from Proposition 1 that $\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)=\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$.
  2. (ii)
    For $j>0$, note that $n\in\{\bar{i}\}$ implies $\gamma_{\text{+}}(\ell_{n})\in\{\gamma_{\text{+}}(\ell_{\bar{i}})\}$, and $\gamma^{\prime}_{\text{+}}(\ell_{n})\in\{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})\}$, i.e.,
$1_{\gamma_{\text{+}}(\ell_{\bar{i}})}(\gamma_{\text{+}}(\ell_{n}))$ $=1,$ (24)
$1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(\gamma^{\prime}_{\text{+}}(\ell_{n}))$ $=1.$ (25)
and hence $\gamma_{\text{+}}(\ell_{n})\notin\{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})\}$, and $\gamma^{\prime}_{\texttt{+}}(\ell_{n})\notin\{\gamma_{\texttt{+}}(\ell_{\bar{i}})\}$, i.e.,
Now, apply Proposition 1 to each of the following cases.
(a) $j=\gamma_{\text{+}}(\ell_{n})$:
$\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)(1-1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j))=\eta_{i}(j)$, using (28); and
(b) $j=\gamma^{\prime}_{\text{+}}(\ell_{n})$:
$\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)(1-1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j))=0$, using (25); and
$\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)(1-1_{\gamma{}_{\text{+}}(\ell_{\bar{i}})}(j))=\eta_{i}(j)$, using (29).
(c) $j\neq\gamma^{\prime}_{\text{+}}(\ell_{n})$ and $j\neq\gamma_{\text{+}}(\ell_{n})$: it follows from (26), (27) that $j\in\{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})\}$ iff $j\in\{\gamma_{\text{+}}(\ell_{\bar{i}})\}$, i.e., $1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j)=1_{\gamma_{\text{+}}(\ell_{\bar{i}})}(j)$. Hence, $\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)(1-1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j))=\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$.
Raising the unnormalized conditionals in (a), (b), and (c) to the power of $\beta>0$, and taking their difference, i.e., $\widetilde{\pi}_{i}^{\beta}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))-\widetilde{\pi}_{i}^{\beta}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$, give the desired result. ∎
VI-B Proof of Proposition 3
Proof.
Recall from (19) that $\rho(i|\gamma_{\text{+}})\propto\frac{\phi_{i}(\gamma_{\text{+}}\!(\ell_{i})|\gamma_{\text{+}}\!(\ell_{\bar{i}}))}{\pi_{i}(\gamma_{\text{+}}\!(\ell_{i})|\gamma_{\text{+}}\!(\ell_{\bar{i}}))}$ and let $\varpi(\gamma_{\text{+}})=\frac{1}{P}\sum_{i=1}^{P}\rho(i|\gamma_{\text{+}})$. Since $\Gamma\!_{\text{+}}$ is a discrete space then $\pi(\gamma_{\text{+}})\in[0,1]$. Furthermore, note from [30] the assumption that $\eta_{i}(j)\in(0,\infty)$ and from (16) it follows for any positive 1-1 $\gamma_{\texttt{+}}\in\Gamma\!_{\text{+}}$ that $\pi(\gamma_{\text{+}})\in(0,1)$. By similar reasoning it follows from (21) for any positive 1-1 $\gamma_{\text{+}}\in\Gamma\!_{\text{+}}$ that $\phi_{i}\in(0,1)$. Hence, $\varpi(\gamma_{\text{+}})\in(0,\infty)$ for any positive 1-1 $\gamma_{\text{+}}\in\Gamma\!_{\text{+}}$.
In [30], it was shown that the standard Gibbs sampler (i.e., SGS) is $\pi$-irreducible. It was also shown in [57] that the TGS extension to generate the Markov Chain $\{\gamma_{\texttt{+}}^{(t)}\}_{t\texttt{=}1}^{\infty}$ is $\pi\varpi$-irreducible. Hence, it follows from [38, Proposition 1] that TGS is reversible with respect to $\pi\varpi$. Furthermore, since the importance weight is $w^{(t)}=\left[\varpi(\gamma_{\texttt{+}}^{(t)})\right]^{\texttt{-}1}$ see [38, (1)],
$\lim_{T\rightarrow\infty}\frac{\sum_{t\texttt{=}1}^{T}w^{(t)}h(\gamma_{\texttt{+}}^{(t)})}{\sum_{t\texttt{=}1}^{T}w^{(t)}}=\sum_{\gamma_{\texttt{+}}\in\Gamma\!_{\text{+}}}\pi(\gamma_{\texttt{+}})h(\gamma_{\texttt{+}}),$
for any bounded test function $h$. The convergence result follows noting that $\left[\varpi(\gamma_{\texttt{+}}^{(t)})\right]^{\texttt{-}1}\!\!\propto P/\left\langle\rho(\cdot|\gamma_{\text{+}}^{(t)}),1\right\rangle$. The stability of the importance weights follows from [38, Proposition 2], i.e., the variance of the weights does not grow with the number of coordinates. Specifically,
$var(w^{(t)})\leq\max_{i,\gamma_{\texttt{+}}^{(t)}}\frac{\pi_{i}(\gamma_{\texttt{+}}^{(t)}\!(\ell_{i})|\gamma_{\texttt{+}}^{(t)}\!(\ell_{\bar{i}}))}{\phi_{i}(\gamma_{\texttt{+}}^{(t)}\!(\ell_{i})|\gamma_{\texttt{+}}^{(t)}\!(\ell_{\bar{i}}))}-1.$